OpenStudy (anonymous):

How do I simplify [(e^x + 1) - (e^x - 1)^-2] + [e^x -1)^-1 e^x] I'll retype it as an equation in a post

OpenStudy (anonymous):

\[e^{x} -(e^{x} -1)^{-2} + (e^{x}-1)^{-1}e ^{x}\]

OpenStudy (anonymous):

\[ = \frac{ -e^x }{ (e^x - 1)^2 } + \frac{ e^x }{ e^x -1}\] Is this so far correct?

OpenStudy (anonymous):

\[= \frac{ -e^x }{ (e^x - 1)^2 } + \frac{ e^x }{ e^x -1} \frac{ (e^x - 1) }{ (e^x - 1) }\]

Directrix (directrix):

The beginning of the problem as first posted does not appear to be the same as the second time you posted the problem. Or, is the second post, the first step you took in solving the problem? If I want to see the original problem, where should I look?

OpenStudy (anonymous):

Thanks. You are correct. I figured out the answer.